1) Suppose economists observe that a $10B increase in government spending raises aggregate demand in the economy by $30B. If this country does not trade and there is no crowding out of private sector spending, what would MPC be equal to?

2) Do you agree or disagree with the following statement? Explain your answer

“An increase in government purchases (G) is usually a more effective policy than an increase in transfer payments when the goal is to eliminate a recessionary gap.”
Answer #1

Question 1:

According to spending multiplier formula:
Change in AD/Change in spending = 1/(1-MPC)
So, 30/10 = 1/(1-MPC)
So, 1-MPC = 1/3
So, MPC = 1-(1/3) = 2/3

Thus, MPC = 2/3 = 0.67

Question 2:

Yes, we agree with the given statement. This can be explained through the multiplier approach.
Government spending multiplier is given by: Change in AD/Change in spending = 1/(1-MPC)
Transfer payments multiplier is given by: Change in AD/Change in transfer payments = MPC/(1-MPC)
MPC lies between 0 and 1. So, 1/(1-MPC) > MPC/(1-MPC). Thus, increase in government purchases would lead to a greater increase in aggregate demand than the same increase in transfer payments. So, increase in government purchases (G) is usually a more effective policy than an increase in transfer payments when the goal is to eliminate a recessionary gap.